MikesArcade.com Coin-op Video Arcade Technical Information Archive
  Home  |  Repair Logs  |  Game Manuals  |  Monitor Manuals  |  Online Store  |  Links  |  Contact Us
Technical Info
  Repair Logs
  Pacman fixes
  Manuals & Schematics
  Monitor Manuals
  Nintendo Game List
  Nintendo Joysticks
  Nintendo Buttons
  Data I/O Promlink 6.10
Spies Wiretap Archives
  Game Conversions
  The Mixed Bag
  PCB Pinouts and DIPs
PDF Game Artwork
  Coin door inserts, etc.
  Instructions cards
  Operation Sheets
  Control Panel & Bezels
Jukebox
  Title Strip Creator
Tutorials
  Component markings...
  Crating a game
  Build your own cabinet
Board Upgrades
  Double Donkey Kong
Wanted
  Items I'm looking for
Miscellaneous
  Links
  Contact Us
MAMETM & MAMEUI
  Visual C Project files
Certifications
  Valid HTML 4.01!
 

Repair Logs


Hippodrome (Data East) PCB repairs

01/13/2000

Tools: Logic probe, Multi-meter, soldering iron, 68k pinouts

Problem: Dead board

This is the same DECO (Data East Corp) CPU board, DE-0297-3 used on Bad Dudes, Dragon Ninja and Robo-Cop. I have a Bad Dudes board set, so I put the ROM board on the known working Bad Dudes CPU board to figure out which 1/2 of the board set was the problem. The ROM board worked just fine on the Bad Dudes board so the problem was isolated to the CPU board. This was an easy way to figure out where to start. :-)

After putting the board back together, I powered up the board and looked for Video sync. It was present on the edge connector. Next checked the reset line on the 68k CPU. It was stuck low. The reset line should be high for normal game operation. I traced the reset line to the 7405 Hex inverter next to the CPU and the output on pin 8, which fed the res,Heline of the CPU was stuck low. It looked like this chip had been replaced before, and soldered straight back to the board. I checked the pin connections for continuity with the multi-meter and discovered pin 5 was not soldered at all to the board. I soldered the pin and reflowed the solder on the other pins. Then I powered on the board. The CPU came to life! I probed the CPU and the address and data lines were pulsing, as they should be. The rest of the board was running also, which was verified by probing several chips and EPROM data lines. But still no picture.

Problem: No Video

I knew I had sync on the edge connector, but re-checked it again. It was still there. I probed the red, green and blue leads on the connector and each one was dead. Nothing! I traced them back on the board, and noticed there was a pretty deep gash in the PCB about 2 inches from the edge connectors. I powered down the board and grabbed my multi-meter. I used the meter to check the connection of the red line at the card edge to FB1 were the trace ran. Open! I did the same with the green and blue to FB2 and FB3 respectively. Open! Looks like the problem. I ran jumper wire for each trace and checked my work with the meter. Looking good! The digital multi-meter happily beeped as I tested the traces. Fired up the board and Viola! Working Hippodrome. A very happy ending and only a couple of hours work. It was interesting to me that all three traces were broken.

Solutions: Fixed the bad solder joints on the 7405 and fixed three traces for red, green and blue.


01/14/2000

Tools: Logic probe, Multi-meter, soldering iron, Razor blade, 68k pinouts

Problem: Dead board

The board was missing one EPROM on the ROM board and one capacitor at NF36 on the main board. After replacing the EPROM and testing the ROM board on the one I fixed above, I knew this ROM board was also good. I installed a 2200pF filter capacitor at NF35. This cap is the strange 3 legged cap and is commonly the first component the trace hits when leaving the JAMMA connector. NF35 just happened to lead to pin C on the JAMMA connector, which is +5. I reconnected the ROM board and powered up. This got the board to at least power up and get the CPU running. But I had nothing but garbage on the screen. I gave up for the night.

After looking at the board on and off over the last day, I noticed the 3 surface mount custom chips had some bent pins with legs smashed together. Using a razor blade, I straightened out the pins and used a magnifying glass to make sure no legs were touching. I reseated the ROM board and applied power. One more working Hippodrome!

Solutions:

  • Replace missing filter capacitor at NF35
  • Program and replace the missing 27C512 EPROM
  • and straighten out the legs on the three custom ICs on the processor board

< Prev Page

 

Home  |  Online Store  |  Repair Logs  |  Manuals and Schematics  |  Monitor Manuals  |  Component Markings  |  Links

All Content Copyright © 2000-2017 by MikesArcade.com